OpenStudy (anonymous):

Solve. 8 - 5(5t - 19) = -72

OpenStudy (anonymous):

8-25t + 95 = -72 -25t +103 = -72 -25t = -72 - 103 -25t = -175 divide both sides by -25 t= -175/ -25 t = 7

OpenStudy (anonymous):

Thanks